In a 2-3 page analysis, answer the following questions Warf Computers, Inc., was founded 15 years ago by Nick Warf, a computer programmer. The small initial investment to start the company was made by Nick and his friends. Over the years, this same group has supplied the limited additional investment needed by the company in the form of both equity and short- and long-term debt. Recently the company has developed a virtual keyboard (VK). The VK uses sophisticated artificial intelligence algorithms that allow the user to speak naturally and have the computer input the text, correct spelling and grammatical errors, and format the document according to preset user guidelines. The VK even suggests alternative phrasing and sentence structure, and it provides detailed stylistic diagnostics. Based on a proprietary, very advanced software/hardware hybrid technology, the system is a full generation beyond what is currently on the market. To introduce the VK, the company will require significant outside investment. Nick has made the decision to seek this outside financing in the form of new equity investments and bank loans. Naturally, new investors and the banks will require a detailed financial analysis. Your employer, Angus Jones & Partners, LLC, has asked you to examine the financial statements provided by Nick. Here are the balance sheets for the two most recent years and the most recent income statement: WARF COMPUTERS Balance Sheet ($ in thousands) 2015 2014 2015 2014 Current assets Current liabilities Cash and equivalents Accounts receivable Inventories Other $ 452 $ 391 668 663 78 Accounts payable $ 519 485 401 716 641 92 Accrued expenses 247 Total current liabilities $ 766 $ 886 Total current assets $1,901 1,800 Long-term liabilities Deferred taxes Long-term debt 330 159 1,148 Fixed assets 1.179 Property, plant, and $4,148 $%3,179 Total long-term liabilities equipment Less accumulated depreciation $1,509 $1,307 1.340 092 Stockholders equity Preferred stock Common stock Capital surplus Accumulated retained $ 21 21 126 779 2.478 ,603 Net property, plant, and equipment Intangible assets and others $2,808 $2,087 709 $3,60 $2,796 126 794 793 Total fixed assets earnings Less treasury stock 126 $3,227 $2,403 192 Total equity Total liabilities and shareholders equity $5,502 $4.596 Total assets $5,502 $4,596Nick has also provided the following information: During the year the company raised $228,000 in new long-term debt and retired $197,000 in long term debt. The company also sold $15,000 in new stock and repurchased $66,000 in stock. The company purchased $1,482,000 in fixed assets and sold $429,000 in fixed assets. WARF COMPUTERS Income Statement ($ in thousands) Sales Cost of goods sold Selling, general, and administrative expense Depreciation Operating income Other income EBIT Interest expense Pretax income Taxes $7,557 4,456 848 248 $2,005 75 $2,080 137 $1,943 776 Current: $605 Deferred 7 Net income Dividends Retained earnings $1,167 $292 $ 875 Angus has asked you to prepare the financial statement of cash flows and the accounting statement of cash flows. He has also asked you to answer the following questions 1. How would you describe Warf Computers cash flows?Dividends Retained earnings $ 292 $ 875 Angus has asked you to prepare the financial statement of cash flows and the accounting statement of cash flows. He has also asked you to answer the following questions 1. How would you describe Warf Computers cash flows? 2. Which cash flow statement more accurately describes the cash flows at the company? 3. In light of your previous answers, comment on Nicks expansion plans. Provide the financial statement of cash flows and the accounting statement of cash flows along with a 2-3 page analysis that thoroughly answers the questions presented with a strong supporting rationale. Your paper should be formatted in APA style with a title page. Due: Sunday by 11:59 p.m., CT Points Possible: 50
Answer #1
PARTICULAR AMOUNT($) AMOUNT($)
Cash at the beginning of the year 391000
OPERATING ACTIVITY:
Net Income 1167000
ADD:Depreciation 248000
:Deferred taxes 171000
ADD:Increase in A/p 34000
Decrease in Inventory 22000
LESS: Increase in other CA 14000
Increase in Accounts Receivables 48000
Decrease in other CL 154000
Net cash from operating Activities 2249000
INVESTING ACTIVITY
sale of fixed assets 429000
purchase of fixed assets (1482000)
Net cash from investment (1053000)
FINANCING ACTIVITY
Equity purchased (66000)
Equity sold 15000
Debt raised 228000
Debt retired (197000)
Dividend Paid (292000)
   increase in long term loan 31000
Net cash from financing (281000)
increase in net cash 915000
cash at the end 1306000

FINANCIAL CASH FLOWS:

cash flow from assets = cash flow paid to creditors + cash flow paid to equity investors

cash flow from assets= Operating cash flow - net capital spending- changes in net working capital

operating cash flow:

operating cash flow= EBIT + depreciation -tax

=2080000+248000-776000=1552000

net capital spending = ending net fixed assets- beginning net fixed assets + depreciation

=3601000-279600+248000=1053000

changes in net working capital= ending NWC- beginning NWC

( CA-CL)-(CA-CL)

(1901000-766000)-(1800000-886000)= 221000

CASH FLOW FROM ASSETS:

PARTICULARS AMOUNT$
operating cash flow 1552000
less: net capital spending (1053000)
changes in NWC (221000)
cash flow from assets 278000

CASH FLOW TO CREDITORS :

PARTICULARS amount
interest paid 137000
less: net borrowing 31000
cash flow to creditors 106000

CASH FLOW TO STOCKHOLDERS:

PARTICULARS AMOUNT
DIVIDENDS PAID 292000
less: equity raised (66000)
debt raised (228000)
cash flow to stockholder (2000)

CASH FLOW TO INVESTORS:

=cash flow to creditors +cash flow to stockholder

=106000+2000=108000

CASH FLOW FROM ASSETS = CASH FLOW TO INVESTORS

Q1 according to warf computer cash flow

*it is having inflows from operating activities whereas outflow from financing activities

*but overall there is cash inflow in the firm which shoe good cash generating ability of the firm

*financial investors can consider investing in this firm because its cash flow are positive and there is less chance of bankruptcy

Q2

from investment point of view, accounting cash flow more accurately describes cash flow at the company as:

* it shows the investors each activity and the cash generation from activities separately

*all expense and revenues to the year , investments and financial position is shown through this statements

*investors can take decision at a glance by seeing the statement that whether they should invest or not

Q3

Nick expansion plan is favorable because:

*even by raising amount of debt, increasing equity and investing in fixed assets it have a positive cash flow according to  accounting cash flow statements